OpenStudy (anonymous):

1/8x squared 4/27x 8/9x Im looking for the LCD of these three fractions

OpenStudy (anonymous):

8 = 2*2*2 27 = 3*3*3 64 = 2*2*2*2*2*2 lcd = 3*3*3*2*2*2*2*2*2 = 27 * 64 = 1728 1728x